Hours 1003

From DIDEAS Wiki
Jump to: navigation, search


  • fix IMU ADCs. Need to sync, implement IIR filter or at least sync the average w/ the 2010 request. need to return the time of the sample.
  • make a renishaw PCA with XLP (ASAP)
  • get to the bottom of the encoder drift (run bench test)
  • program to verify LTF operation (automated)
  • teach CA to run the benchtest.
  • fabricate more SWIFI PCBs
  • fix wifi_fast text log file precision
  • wifi fast to get precision from the vlist file.
  • wifi_fast to have a "duration" that causes it to stop periodically.
  • LTF2 experiment using JR3
    • write C code for same
  • close loop control of output torque
  • code to push SVN between servers.
  • get DD-WRT VPN operational (solves subnet problem)
  • improve the python paths.

RIC todo

  • update SWIFI firmware on S4.
  • email calb data to RIC
  • UDP messages to control WIFI fast.

lab todo

  • code to push SVN changes from trunk into each users space.

MAR 2010

mon mar 1

  • 30 mins emails re firmware and config w/GG, multiple LTF plan w/GG
  • admin for severl hrs.
  • 3:45pm - 6:45p
    • discuss w/DG LTF2, and moving to back.
    • work w/CA to get LTF2 back on network
    • discuss new network architecture for swifis (lab net) and BSt. w/ED
    • email w/GG re building
    • 9 phone calls from 3:20p - 8:20p

  • 1.5 hrs.
    • drive to microcenter.
    • experiment with new access point
    • setup dd-wrt

tue mar 2

  • 2:30pm - 6:50p
  • continue access point config

setup PC at bimneny

  • setup LTF, get views working
  • discuss network and AP w/ED
  • 14 phone calls w/IW from 1:17-6:40p, 9:45-10:30p


  • investigate LTF network issue
  • setup VPN at Binney

wed mar 3

  • 10:50a - 6:10p
    • 6 IW calls
    • work with LTF1 and DG to determine cause of excessive temperature rise
    • cause traced to apparent encoder drift.

thur mar 4

  • 8.45a - 11:40a
    • experiments with LTF1 to confirm encoder drift.
    • modify plot_ltf to support viewing of new experiment data.
    • email DG re LTF1 status and plan
  • 4:50p - 9:10p, -25m
    • reviewed status of WIP4d for submission
    • discuss w/GG MIT ankle requirements.
    • work w/DG to diagnose LTF1 motor encoder failure
    • improve plot_ltf

fri mar 5

  • 8:30a - 1pm (-30m)
    • restart LTF1, study the cause of the stoppage, investigate RLFv2 for untrapped errors and fix.
    • coordination with DG re LTF
    • update RC w/status of LTF and next week.
    • email to RC, HH re situation

sun mar 7

  • 25 min, phone w/ED

mon mar 8

  • 1:05p-2:50p
    • prepare AP DD-WRT for lab, test w/S5
  • 6p-7p
    • discuss robot status with GG
    • test pser2h-b, find component not installed, CA installs and SN #1-3 tested successfully, to be delivered to MIT.
  • 9:15p-10:40p
    • phone ED re issues with LTF1
    • update ip devices wiki w/ Binney network info.
    • test communications with S3 and L5
    • debug LTF1, assign network device address, get vncserver working.
    • go to Binney to looking to problems with LTF1, insert JR3 cable, still no comedi device

tue mar 9

  • 8am-
    • phone w/GG
    • review GG robot torque / beta code
    • bring up LTF1, problems with running hot and lack of thrust
  • -3:20p

wed mar 10

  • 9a-2:00p -30min
  • 9 calls, 9:38a-3pm
    • review LTF1 (for past 24 hrs), save plots for encoder drift. need to investigate this encoder.
    • add ability for udp_matlab to sent to multiple IPs. Add support for same to wifi_fast.py
    • get JR3 working with RLFv2, still crashes and times out. Add log file capability
    • discuss LTF2 status with DG
    • dealing with IMU connector orientation. Rev200 has a BOM ECO (undocumented) - now using RA header.
  • phone w/ED 30min @ 10pm
  • 1 hr
    • experiments with DD-WRT to support VPN into the router.

thur mar 11

  • 7:50a - 2:30p (-30 min) (-30min)
    • improve wifi_fast plotter to always read the full UDP buffer, then uses variable plotter frame rate.
    • test with plot_lifefix
    • temporary fix to wifi_fast fancy_list_output
    • re-arrange python support scripts, improve python_paths registry, start merge of LTF and PC SVN dirs.
  • 30 mins :
    • review LTF and restart
    • review issues with signals, try to determine cause

fri mar 12

  • 8:30a-12:45p
    • mcmaster order for damping foam and tools
    • review LTF and debug signals.
    • phone w/ED
    • SVN distro
  • 12:30p-2:050p
    • try to get LTF2 operational
    • phone w/RC

sat mar 13

  • worked most of the day.
    • LTF1 not running, research the cause. need to decode data files
    • hall sensor study, find 0.5 rad/sec of effect from motor magnets
    • email to RC on how we need to proceed.
    • rewrite hall sensor code to make calb table ready

sun mar 14

  • 8:30p - 10:30p +30.
    • commit and document SVN changes.
    • experiment with hall sensor used in feed back loop.

mon mar 15

  • 11:45a-2:35p
    • email RC re hall sensor
    • review IMU changes from Rev100-Rev200, respond to MR email on same and ECOs.
    • write test code for torque estimation, email to ZH and GG
    • phone w/RD, ED, GG, DG

  • 4p-7:15p
    • robot joint torque code development
    • experiments with RTW compilation
    • verify LTF1 operation, encoder is connected to wrong port, started LTf1, discuss status w/DG
  • 9p-10:15p
    • debug,verify operation of torque sensor

tue mar 16

  • 10:50a- 1:50p
    • review LTF1 operational status
    • investigate garbage output associated with "ACO" debug, was printf %d again
    • change S3 build to be -Os
  • 2:00p - 3:00p
    • phone w/RC
  • 3:30p-8:30p
    • phone w/ED
    • discuss torque and hall calibration with RC, then IMU with RC and ZB
    • discuss w/ED python wrapper for copley and methods of capturing PDOs
    • assist ZH to re-install python to get numpy to work on laptom for IMU collection
    • Q&D test with renishaw encoder and EC30 - there is interference, discuss with CW
    • meet with RC to discuss uses of renishaw, how to get velocity. Reviewed how to use the spare port on the SC/MC. There are 2 (SC / MC) analog inputs that could be repurposed as digital IOs. There is also a MC digital IO
    • review IMU source code, discuss with ZH and RC
    • discuss wire and having cables made with CA
    • work to bring up LTF2 - source code needs to be organized
  • 9:30p - 10:45p
    • re-work source code on LTF. not sure why JR3 cant be open when using source files on PC

wed mar 17

  • 9:00a-2:10pm -1.15hr
    • review LTF1, diagnose cause of failure, email DG
    • setup new XP box, LTF viewer PC
    • work to resolve wifi_fast issues.
    • DG fixes LTF1, I restart and still doesn't work.
    • review MC source for how to write data structures w/mem write.
  • new firmware - 2 hrs - verified message routing by unit ID
  • installed DD WRT on 54GL - VPN now works.

thur mar 18

  • 30 mins - phone calls from GG re SM rebuild of SWIFI
  • 40 mins discuss torque calibration w/MikeE, delivery of testjig, and control methods.
  • new firmware - 4.5 hrs - generate DMA based message transmission firmware
  • new firmware - 2 hrs - verify operation of transmission control object.

fri mar 19

  • review LTF1 status
  • 3 hr - new firmware development, SPI interface development

sat mar 20

  • review LTF1 status

sun mar 21

  • review LTF1 status

mon mar 22

  • review LTF1 status
  • 2 hrs w/RC discuss pyramid and hall calibration, S3 status, and renasha

tue mar 23

  • 1:25p-4:50p phone
  • 15 mins - discuss LTF with DG
  • 20 mins - discuss Copley wrapper w/ED
  • 20 mins - discuss LTF logon / control training w/ED
  • 15 mins - deal with S4 SWIFI firmware bug.
  • 30 mins - discuss LTF access wi DG/ED
  • 1:20 - meet w/ED and DG to discuss LTF and how to setup
  • 15 min - set up XP box, pick up S3

wed mar 24

  • 50 mins - deal with LTF1 problem. send email on same.

thur mar 25

  • NA

fri mar 26

  • 10:15a - 1:30p
    • review calibration data for S3, S5. emails to ZH and RC
    • create wiki page for Pser2h_revb
    • email CA, RC re SWIFI,encoder issue.
  • 4:20p - 9:35p
    • discuss calibration poly with ZH
    • fix issues with python / benchtest
    • discuss SWIFI w/GG
    • experiment with renishaw encoder.

sat mar 27

  • 1.5 hrs - Renishaw PCB design

sun mar 28

  • 1 hr Renishaw PCB design
  • 1pm - 7:20p
    • relocate LTF viewer PC to robot room
    • plot LTF improvements
    • investigate LTF2 status, repair encoder, verify operation
    • setup calbstand to be able to reprogram LTF2
    • MC firmware mods to more cleanly use the JR3 over serial
    • testing of program that sends JR3 as q15.
  • 7:55p - 10p at lab

mon mar 29

  • renishaw PCB design
  • discuss status w/RC

tue mar 30

  • renishaw PCB design
  • 15 min IW phone
  • dinner w/RC

wed mar 31

  • 10min w/RC
  • 15min w/IW